# Note on Min-$k$-Planar Drawings of Graphs

### v4

@@ 4 Conclusions
Many papers in the graph drawing area deal with only simple drawings, either as a convenient simplification of the general case, or as a strict condition in the definition. However, some works are not clear in distinguishing between the two situations and this can bring troublesome problems. In this regard, we would like to mention, for instance, a similar past confusion thoroughly studied in the remarkable paper of Pach and Tóth [DBLP:journals/jct/PachT00] (entitled “Which crossing number is it anyway?”).
We have demonstrated that a careful distinction (between simple / non-simple drawings) is surely necessary when considering the recent min-$k$-planar graphs. Our note brings a natural open question about which of the published results of [DBLP:conf/gd/BinucciBBDDHKLMT23] concerning simply min-$k$-planar graphs with $k\geq 2$ remain valid also for general min-$k$-planar graphs.

### v5

@@ 4 Conclusions
Many papers in the graph drawing area deal with only simple drawings, either as a convenient simplification of the general case, or as a strict condition in the definition. However, some works are not clear in distinguishing between the two situations and this can bring troublesome problems in the future. In this regard, we would like to mention, for instance, a similar past confusion thoroughly studied in the remarkable paper of Pach and Tóth [15] (entitled “Which crossing number is it anyway?”).
We have demonstrated that a careful distinction (between simple / non-simple drawings) is surely necessary when considering the recent min-$k$-planar graphs. Our note brings a natural open question about which of the published results of [4] concerning simply min-$k$-planar graphs with $k\geq 2$ remain valid also for general min-$k$-planar graphs.